OpenStudy (anonymous):

Point E is located at (–2, 2) and point F is located at (4, –6). What is the distance between points E and F?

OpenStudy (anonymous):

\[\sqrt{52}, \sqrt{28}, 10, \sqrt{20}\]

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

\[d=\sqrt({{-2_{1}}}-4_{2})^{2}+(2_{1}+-6_{2})_{2}^{?}\] Is this right?

hartnn (hartnn):

don't write the subscripts again, \(d = \sqrt {(-2-4)^2+(2-(-6))^2}\) can you simplify ?

OpenStudy (anonymous):

oh ok so when i do the math I get 70 as d. Is that correct?

OpenStudy (anonymous):

I feel as if I did something wrong though...

hartnn (hartnn):

6^2 is 36 8^2 is 64 and whats 36 + 64 ?

OpenStudy (anonymous):

100

hartnn (hartnn):

sqrt (100) =.. ?

hartnn (hartnn):

did you get how it is 100 ?

OpenStudy (anonymous):

Ohhhh 10. now I get it. xD

hartnn (hartnn):

cool :) yes, distance = 10 :)

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

welcome ^_^